[Hallc_running] [New Logentry] Update on Hall-C Running

latif at jlab.org latif at jlab.org
Sat Jan 27 12:15:02 EST 2018

Logentry Text:
The beam has been comparatively stable and we are making progress since yesterday day-shift. We are taking CT production data for the lowest Q^2 point (8.01 GeV^2) at three pass. This data will confirm previous measurement. Our goal is to take 36 hours of data. So far we have taken around 12 hours of data (with carbon target). Major downtime since yesterday's update includes -- 1) HMS dipole and magnets tripped around 5:00 PM yesterday for lack of He flow which were restored around 11:50 PM.  2) The accelerator experienced 3 pass separator trip this morning around 6:00 AM and was restored around 8:00 AM.

There will be no RC meeting today, however we plan to meet in the CH conference room tomorrow at 1:30 PM to give an update on weekend running. The meeting will be available through bluejeans meeting ID: 543 903 672.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3521986
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180127/0bd2ff95/attachment.html>

More information about the Hallc_running mailing list